OpenStudy (anonymous):

Can someone do this step by step please? how do i integrate this: π∫(tanx)^(2)dx going from x=0 to x=π/4

OpenStudy (anonymous):

it's okay if you don't solve it all the way. i just can't think of the aniderivative of tanx^2

OpenStudy (anonymous):

\[\pi \int\limits_{0}^{\pi/4}(\tan ^{2}x)dx \] integrate and you will obtain \[\pi((\tan x)-x) _{0}^{\pi/4}\] then solve the equation.

OpenStudy (anonymous):

the antiderivative is tanx-x?

OpenStudy (anonymous):

yup.

OpenStudy (anonymous):

okay thank you! can you help me with my question before this one? it's just finding the antiderivative

OpenStudy (anonymous):

can you post the question back here..

OpenStudy (anonymous):

0 tmmss5533 Can someone do this step by step please? how do i integrate this: π∫e^(4)- e^(4x^(x))dx going from x=0 to x=1

OpenStudy (anonymous):

just finding the antiderivative is fine

OpenStudy (anonymous):

That integration is a little bit complicated. you need to use integration by parts as the function itself is rather a complex function.

OpenStudy (anonymous):

e^(4x^(2)) does that help a little?

OpenStudy (anonymous):

sorry i typed it wrong >.<